# Half-integer thermal conductance in the absence of Majorana mode

### v1

@@ Abstract
Considering a range of candidate quantum phases of matter, half-integer thermal conductance ($\kappa_{\text{th}}$) is believed to be an unambiguous evidence of non-Abelian states. It has been long known that such half-integer values arise due to the presence of Majorana edge modes, representing a significant step towards topological quantum computing platforms. Here we break this long-standing paradigm, reporting a comprehensive theoretical and experimental study where half-integer two-terminal thermal conductance plateau is realized employing Abelian phases. Our proposed setup features a confined geometry of bilayer graphene, interfacing distinct particle-like and hole-like integer quantum Hall states. Each segment of the device exhibits full charge and thermal equilibration. Our approach is amenable to generalization to other quantum Hall platforms, and may give rise to other values of fractional (electrical and thermal) quantized transport. Our study demonstrates that the observation of robust non-integer values of thermal conductance can arise as a manifestation of mundane equilibration dynamics as opposed to underlying non-trivial topology.
@@
In this work, we demonstrate that half-integer thermal conductance can be realized without the presence of a Majorana mode. We utilize n-p-n heterojunctions consist of symmetry broken $\nu^{\prime}=-1$ region sandwiched between two $\nu=2$ integer QH state in the zeroth Landau level (ZLL) of bilayer graphene (BLG) to artificially create a $\frac{1}{2}G_{0}$ electrical conductance state, where $G_{0}=e^{2}/h$, is the electrical conductance quanta. The high tunability of symmetry-broken flavors (spin, valley, and orbital)[25, 26, 27, 28] in the ZLL of BLG — through both magnetic and electric fields — enables full equilibration between the co-propagating electron-hole (e-h) modes for which the two-terminal electrical conductance becomes, $G_{\nu-\nu^{\prime}-\nu}=\frac{|\nu||\nu^{\prime}|}{|\nu|+2|\nu^{\prime}|}G_{0}=\frac{1}{2}G_{0}$[17] (see "Methods" and supplementary). To test the idea that, two-terminal thermal conductance of such QH n-p-n junction also leads to a value of $\frac{1}{2}\kappa_{0}T$ (which is proposed by our theory as discussed in method and Supplementary Information (SI) section 12), we employ a unique device geometry with three arms as shown schematically in Fig. 1a. This geometry enables the creation of a temperature difference ($\Delta T$) across the n-p-n junction while maintaining zero net charge current, despite heat transport via chaotic (e-h) mixing across the junction. Our data suggests, full equilibration of both charge and energy leading to thermal conductance of $\sim\frac{1}{2}\kappa_{0}T$, consistent with our theoretical calculations. This challenges our current understanding as half-quantization of heat conductance is a unique signature of a Majorana mode, which makes this work is of fundamental importance. Our results also confirm that the Wiedemann-Franz law remains valid for engineered fractional values of thermal conductance, in contrast to the behavior observed in anyonic heat flow [11, 5, 13, 14]. We reproduce this result in another device consists of a p-n-p junction with $\nu=-2$ and $\nu^{\prime}=1$, reinforcing the reliability of our findings.
@@ Results:
Our work puts on the table the issue of robustness of quantum transport platforms to external perturbations. One may compare topology-based resilience to resilience underlined by strong equilibration dynamics. Such a study can use case studies of other equilibration-engineered fractional values of thermal and electrical conductance, as well as also serving a test-bed for the validity of the Wiedemann-Franz law. This might be particularly interesting when implemented to platforms which (in the absence of equilibration) exhibit charge fractionalizaton giving rise, e.g., to charged and neutral eigenmodes [42, 43, 44, 45]. Generalizations to fractional quantum Hall boundaries (e.g., with charge and neutral modes [46, 44] or charge and spin modes [47]) should follow. This work also opens up an intriguing direction of incorporating equilibration dynamics (e.g. [48]) into the physics of dilute colliding beams[49, 50, 51, 52], involving both fermions and anyons.
@@ Methods
Acknowledgements
U. R. and S. M. thanks Arup Kumar Paul for useful discussions. We thank the International Centre for Theoretical Sciences (ICTS) for participating in the program - Condensed Matter meets Quantum Information (code: ICTS/COMQUI2023/9), where the collaboration was initiated. S.M. was supported by the Weizmann Institute of Science, Israel Deans fellowship through Feinberg Graduate School. Y. G. was supported by the InfoSys Chair, IISc, Bangalore. S.M. and Y. G. were also supported by the Minerva Foundation and grant no 2022391 from the United States–Israel Binational Science Foundation (BSF), Jerusalem, Israel. Ankur Das was supported by IISER, Tirupati Startup grant and ANRF/ECRG/2024/001172/PMS. M.G. has been supported by the Israel Science Foundation (ISF) and the Directorate for Defense Research and Development (DDR&D) through Grant No. 3427/21, the ISF Grant No. 1113/23, and the US-Israel Binational Science Foundation (BSF) through Grant No. 2020072. A.D. thanks the Department of Science and Technology (DST) and Science and Engineering Research Board (SERB), India, for financial support (SP/SERB-22-0387), (DST/NM/TUE/QM-5/2019), and also acknowledges funding through the Intensification of Research in High Priority Areas programme of the Science and Engineering Research Board (Grant No. IPA/2020/000034). A.D. also thanks CEFIPRA project SP/IFCP-22-0005. Growing the hBN crystals received support from the Japan Society for the Promotion of Science (KAKENHI grant nos. 19H05790, 20H00354 and 21H05233) to K.W. and T.T. The authors gratefully acknowledge the use of Blender (https://www.blender.org/) and Inkscape (https://inkscape.org/) — both free and open- source software —- for the creation and refinement of figures presented in this work.
Competing interests
The authors declare no competing interests.
See pages 1-18 of SI.pdf

### v2

@@ Abstract
Considering a range of candidate quantum phases of matter, half-integer thermal conductance ($\kappa_{\text{th}}$) is believed to be an unambiguous evidence of non-Abelian states. It has been long known that such half-integer values arise due to the presence of Majorana edge modes, representing a significant step towards topological quantum computing platforms. Here, we challenge this prevailing notion by presenting a comprehensive theoretical and experimental study where half-integer two-terminal thermal conductance plateau is realized employing Abelian phases. Our proposed setup features a confined geometry of bilayer graphene, interfacing distinct particle-like and hole-like integer quantum Hall states. Each segment of the device exhibits full charge and thermal equilibration. Our approach is amenable to generalization to other quantum Hall platforms, and may give rise to other values of fractional (electrical and thermal) quantized transport. Our study demonstrates that the observation of robust non-integer values of thermal conductance can arise as a manifestation of mundane equilibration dynamics as opposed to underlying non-trivial topology.
@@
In this work, we demonstrate that a robust half-integer thermal conductance can be realized in a complex structure, which is not subject to fine-tuning of the system’s parameters. Such half-integer values are not a manifestation of an underlying non-Abelian topology, and in particular they do not represent an underlying Majorana mode. We utilize n-p-n heterojunctions consist of symmetry broken $\nu^{\prime}=-1$ region sandwiched between two $\nu=2$ integer QH state in the zeroth Landau level (ZLL) of bilayer graphene (BLG) to artificially create a $\frac{1}{2}G_{0}$ electrical conductance state, where $G_{0}=e^{2}/h$, is the electrical conductance quanta. The high tunability of symmetry-broken flavors (spin, valley, and orbital)[25, 26, 27, 28] in the ZLL of BLG — through both magnetic and electric fields — enables full equilibration between the co-propagating electron-hole (e-h) modes for which the two-terminal electrical conductance becomes, $G_{\nu-\nu^{\prime}-\nu}=\frac{|\nu||\nu^{\prime}|}{|\nu|+2|\nu^{\prime}|}G_{0}=\frac{1}{2}G_{0}$[17] (see "Methods" and supplementary). To test the idea that, two-terminal thermal conductance of such QH n-p-n junction also leads to a value of $\frac{1}{2}\kappa_{0}T$ (which is proposed by our theory as discussed in method and Supplementary Information (SI) section 12), we employ a unique device geometry with three arms as shown schematically in Fig. 1a. This geometry enables the creation of a temperature difference ($\Delta T$) across the n-p-n junction while maintaining zero net charge current, despite heat transport via chaotic (e-h) mixing across the junction. Our data suggests, full equilibration of both charge and energy leading to thermal conductance of $\sim\frac{1}{2}\kappa_{0}T$, consistent with our theoretical calculations. This challenges our current understanding as half-quantization of heat conductance is a unique signature of a Majorana mode, which makes this work is of fundamental importance. Our results also confirm that the Wiedemann-Franz law remains valid for engineered fractional values of thermal conductance, in contrast to the behavior observed in anyonic heat flow [11, 5, 13, 14]. We reproduce this result in another device consists of a p-n-p junction with $\nu=-2$ and $\nu^{\prime}=1$, reinforcing the reliability of our findings.
@@ Results:
Our work puts on the table the issue of robustness of quantum transport platforms to external perturbations. One may compare topology-based resilience to resilience underlined by strong equilibration dynamics. Such a study can use case studies of other equilibration-engineered fractional values of thermal and electrical conductance, as well as also serving a test-bed for the validity of the Wiedemann-Franz law. This might be particularly interesting when implemented to platforms which (in the absence of equilibration) exhibit charge fractionalizaton giving rise, e.g., to charged and neutral eigenmodes [42, 43, 44, 45]. Generalizations to fractional quantum Hall boundaries (e.g., with charge and neutral modes [46, 44] or charge and spin modes [47]) should follow. This work also opens up an intriguing direction of incorporating equilibration dynamics (e.g. [48]) into the physics of dilute colliding beams[49, 50, 51, 52], involving both fermions and anyons. We also note that half-integer thermal conductance that does not involve non-Abelian modes may appear in other contexts[53].